exp. the fomulaW= 1/2.stress.strain

Grade:10

1 Answers

chadalawada sai krishna
12 years ago

here, W=elastic potential energy per unit volume

elastic potential energy(U)=(Youngs modulous of wire)(area of the wire)(change in length of the wire)2/twice the original length of the wire                       =YA(Δl)2/2l

threrefore,W=U/volume

                  =YA(Δl)2/2Al2

                      =(Δl/l)(y.Δl/2l)

                  =1/2.(strain)(y.strain)

                  =1/2.(stress)(strain)